# Magnetic Flux Emergence in Binary Neutron Star Remnants

### v1

@@ III Results
Figure 1: A slice of the $r$-$\phi$ plane at $t=0.39~{}\textrm{ms}$ comparing $|B|$ from H1e16_5 (left) to the tracer scalar $s_{1}$ from H0_5 (right). The dashed black line marks the surface of the star, and the dot-dashed green lines mark the boundaries of the initial position of the flux tube. The nearly identical evolution indicates that the dynamics of H1e16_5 are driven primarily by hydrodynamic buoyancy from the original perturbation rather than magnetic buoyancy.
@@ III.1 Effects of magnetic field strength
It is easy to verify in the Newtonian limit that this reduces to the usual Newtonian expression for the total non-gravitational energy in MHD. This expression has contributions from the internal, kinetic, and magnetic energy of the fluid. Similar to Ref. [38], we express the kinetic energy $E_{K}$ and magnetic energy $E_{B}$ as
@@
where $dV$ includes the appropriate volume element. Since $\tau$ does not include the gravitational energy of the system, it is not conserved. However, $-\alpha(T^{0}_{0}+\rho u^{0})$ does include these contributions and is therefore conserved in stationary spacetimes. It is easy to verify that $\alpha T^{0}_{0}+\alpha\rho u^{0}+\tau\approx-\rho\Phi$ in the Newtonian limit, where $\Phi$ is the gravitational potential. We therefore define
@@
which we use to normalize the energy of our system. We caution that $E_{\mathrm{tot}}$ is not precisely invariant in due to boundary effects and truncation error from evolving $\tau$ instead of $\alpha(T^{0}_{0}+\rho u^{0})$. However, we have confirmed that our results are the same if we normalize by $E(t=0)$ or $E(t)$, indicating that conservation violations are insignificant for our analysis.
@@
Fig. 2 corroborates this picture; for H1e17_5, $E_{K}$ peaks at a lower energy and begins to decay sooner, indicating that this case is also stable. The magnetic energy decays as well, albeit more slowly than the kinetic energy. If we consider a volume $V=A\ell$, where $A$ is some changing cross-sectional area and $\ell$ is an approximately fixed length, flux conservation implies that $E_{B}\approx B^{2}V\propto A^{-1}$, so we attribute this slow decay in magnetic energy to the slow diffusion or expansion of the flux tube as demonstrated in Fig. 3.
@@
For H7e17_5, which has the strongest magnetic field, the flux tube successfully emerges. This emergence process is shown in Figs. 4 and 5. The flux tube stretches out into a characteristic umbrella shape due to the decreasing scale height as it rises, similar to what occurs in solar studies (e.g., Fig. 4 of Ref. [39]). As it reaches the surface, the flux tube rapidly expands and pushes out a wave of material in front of it (see Fig. 6).
@@
Figure 6: A comparison of $|B|$ (top) and $\rho$ (bottom) in the $r$-$\phi$ plane from H7e17_5 at $t=0.34~{}\textrm{ms}$. Much of the flux tube remains near the surface, but the wave of material pushed out by it is also visible. The dashed black line marks the location of the stellar surface from the TOV solution, and the dot-dashed green lines mark the boundaries of the flux tube at $t=0$.
To measure the emergence time and properties of the entrained material, we integrate the mass flux $\dot{M}$ across surfaces at $r=8.35~{}\textrm{M}_{\odot}$, $r=8.4~{}\textrm{M}_{\odot}$, $r=8.5~{}\textrm{M}_{\odot}$, $r=8.6~{}\textrm{M}_{\odot}$, and $r=8.7~{}\textrm{M}_{\odot}$. We define the emergence time as the peak mass flux $\dot{M}$ across the surface at $r=8.35~{}\textrm{M}_{\odot}$, which is the approximate location of the remnant’s surface. By looking for where $\dot{M}$ peaks in the other integration surfaces, we can measure the approximate speed of the flux tube at emergence, which we do by fitting a quadratic polynomial for $r(t)$ to these peak times and taking its time derivative.
Fig. 7 shows the time-integrated mass across each integration surface along with the estimated emergence time. The flux tube emerges at $t_{\mathrm{em}}\approx 0.216~{}\textrm{ms}$ for H7e17_5 with a speed of $v_{\mathrm{em}}\approx 0.038c$. The total entrained mass peaks around $M\approx 1.5\times 10^{-3}~{}\textrm{M}_{\odot}$ shortly after emergence, but it rapidly declines as material falls back onto the remnant. The mass across each integration surface also rapidly decreases with distance, suggesting that little to no mass is unbound [^1] A better way to estimate unbound ejecta is using the geodesic and/or Bernoulli criteria, but this must be done farther away from the remnant to be accurate. Because our simulations only span $[-\pi/9,\pi/9]$ in the $\theta$ direction with periodic boundary conditions, the emerging fluid begins to interact with itself not long after passing the surface at $r=8.7~{}\textrm{M}_{\odot}$. Therefore, the dynamics of the fluid will not be reliable at a radius where a proper ejecta calculation would be meaningful..
@@
Unsurprisingly, Fig. 2 demonstrates that E7e17_5 has larger $E_{B}$ and $E_{K}$ than in the weak-field runs. $E_{K}$ follows the same qualitative trend as the other runs, with a rapid rise followed by a peak and slow decay in time (possibly with small oscillations). Unlike the other runs, however, $E_{B}$ immediately decays, while it remains nearly constant in E1e17_5 and grows for some time in both E1e16_5 and E6e15_5. There is also a crossover between $E_{B}$ and $E_{K}$ at the same approximate time emergence occurs, though, as will be discussed later, this may simply be coincidence.
@@ III.2 Effects of density perturbation
We now consider the effects of the perturbation by analyzing the non-perturbed cases, beginning with E1e17_0. Since E1e17_5, which has a density perturbation, does not emerge, it is not surprising that E1e17_0 does not, either, and instead rises only for a short time before stalling. As shown in Fig. 8, $E_{K}$ peaks at ${\sim}1/3$ its value in E1e17_5, and it rapidly decays by nearly an order of magnitude in ${\sim}0.3~{}\textrm{ms}$. Additionally, though the evolution of $E_{B}$ in E1e17_0 is initially similar to E1e17_5, it does not experience the same slow rise, but instead slowly decays over time.
Figure 8: Like Fig. 2, but comparing the cases with no perturbation to their perturbed counterparts. The dotted vertical lines mark the emergence times for H7e17_0 and H7e17_5.
Comparing the E7e17_0 case to E7e17_5, we find that the perturbation has very little effect on the dynamics of the system. This is consistent with our initial conditions, as the change in density from our equilibrium conditions on the flux tube ($\Delta\rho/\rho_{0}\approx 0.6$) is much larger than the change in density from the perturbation ($A=0.05$). According to Fig. 8, the presence of a perturbation does not affect the energetics of the system in any significant way. However, there is a small effect on both the emergence time and the mass pushed out by the emerging flux tube. Fig. 7 shows that emergence in E7e17_0 is slightly delayed to $t_{\mathrm{em}}\approx 0.224$ and slightly slower at $v_{\mathrm{em}}\approx 0.034c$ compared to E7e17_5, though the entrained mass is slightly larger. Nevertheless, the mass again rapidly decays with distance, suggesting that the amount of unbound ejecta is small.
@@ III.4 Effects of resolution
There is little consistency in the merger time, and the effects of surface oscillations make it difficult to determine exactly where emergence occurs, particularly for the LR case. Though the flux tube is well resolved for the LR, MR, and HR cases (and arguably adequately resolved for the VLR case), it is likely that we are not in a completely convergent regime, even if we ignore the measured emergence times.
@@ IV Discussion
For our magnetic field configuration, we observe that only extremely large fields are capable of emerging. Because our field is a single twisted loop surrounding the entirety of the star, however, there is a strong magnetic tension force which, using Eq. 24, initially exceeded $50\%$ of the gravitational force on the flux tube in some of our tests. Alternate field configurations may be less susceptible to this tension force, but numerical simulations suggest that post-merger fields, independent of their initial (realistic) conditions, are dominated by a large-scale toroidal component [14, 40], which limits the possibilities for alternate field structures.
@@
Even if this were not the case, this likely is only important for fields of ${\sim}10^{17}~{}\textrm{G}$; we observe that the evolution of the flux tube at weaker field strengths around $10^{16}~{}\textrm{G}$ are dominated by ordinary hydrodynamic buoyancy. Therefore, it seems unlikely that altering the field configuration is enough to help these features to emerge. We do not consider the effects of rotation in our model, but we remind the reader that rotation tends to inhibit convective processes, not enhance them [41, 42, 12].
@@
This result may initially seem at odds with works suggesting that strong magnetically driven outflows may develop from a BNS remnant, such as Most and Quataert [9] and Combi and Siegel [10]. This is not necessarily the case; the demonstrated flare-like features and outflows in Most and Quataert [9] originate from strong magnetic fields exceeding $10^{17}~{}\textrm{G}$. Similarly, the post-merger jet in Combi and Siegel [10] develops in a remnant with an average toroidal field strength nearing and isolated pockets exceeding $10^{17}~{}\textrm{G}$. However, we again stress that such field strengths are atypically large compared to observation of magnetars and what is expected in realistic merger remnants. Even under these circumstances, such fields are not a guarantee that magnetically-driven outflows develop from the remnant. Musolino et al. [43], for example, suggest that, even with an extremely large magnetic field, these outflows may instead originate from the disk.
@@
Similar to alternate configurations, buoyant instabilities like the Parker instability are likely only important for a small range of field strengths in excess of $10^{17}~{}\textrm{G}$. Despite a strong longitudinal perturbation, the Parker instability does not appear to develop in H1e17_5; rather, the flux tube oscillates. This agrees with results from Ref. [43], for instance, which suggest that the post-merger disk can be unstable to the Parker instability, but the remnant itself should be stable, even for a very large magnetic field. Larger fields like H7e17_5 may be Parker-unstable, but the emergence time scale is so short that the instability cannot develop quickly enough to be dynamically important (as seen when compared with H7e17_0). There likely exists some middle ground between these two states where the Parker instability develops quickly enough to accelerate the emergence time scale, but the required field strengths may not be achievable in a realistic merger setting.
@@
As we consider cases which do emerge, it is interesting to note that whether or not emergence occurs seems to depend only mildly on resolution so long as the buoyant feature is adequately resolved. Though this is a necessary condition for emergence, one would not necessarily expect it to be sufficient; under-resolving $H_{\rho}$ could lead to an incorrect density difference between the ambient medium and the interior of the buoyant feature, therefore leading to an incorrect emergence time scale. Similarly, under-resolving $H_{P}$ could cause the buoyant feature to expand too much or too little as it rises, thus indirectly affecting the density difference and limiting buoyancy. In our runs, $H_{\rho}\approx 2.43~{}\mathrm{km}$ and $H_{P}\approx 827~{}\mathrm{km}$ at the shear layer, but both these quantities will rapidly decrease as the flux tube approaches the surface. Therefore, $H_{\rho}$ will only be moderately resolved at the beginning for the VLR run, and $H_{P}$ will only be resolved by the HR and possibly MR runs to start. Though we see some differences in the emergence time (which may not be entirely accurate for H7e17_5_B), they are small enough that they would be insignificant in the context of a full merger simulation. This suggests that the timescales shown in global BNS runs with $\Delta x\lesssim 180~{}\mathrm{m}$ are likely reasonable despite the limited resolution.

### v2

@@ III Results
Figure 1: A slice of the $r$-$\phi$ plane at $t=0.39~\textrm{ms}$ comparing $|B|$ from H1e16_5 (left) to the tracer scalar $X$ from H0_5 (right). The dashed black line marks the surface of the star, and the dot-dashed green lines mark the boundaries of the initial position of the flux tube. The nearly identical evolution indicates that the dynamics of H1e16_5 are driven primarily by hydrodynamic buoyancy from the original perturbation rather than magnetic buoyancy.
@@ III.1 Effects of magnetic field strength
It is easy to verify in the Newtonian limit that this reduces to the usual Newtonian expression for the total non-gravitational energy in MHD. This expression has contributions from the internal, kinetic, and magnetic energy of the fluid. Similar to Ref. Cook et al. (2025b), we express the kinetic energy $E_{K}$ and magnetic energy $E_{B}$ as
@@
where $dV$ includes the appropriate volume element. Since $\tau$ does not include the gravitational energy of the system, it is not conserved. However, $\alpha(T^{0}_{0}+\rho u^{0})$ does include these contributions and is therefore conserved in stationary spacetimes. It is easy to verify that $\alpha T^{0}_{0}+\alpha\rho u^{0}+\tau\approx-\rho\Phi$ in the Newtonian limit, where $\Phi$ is the gravitational potential. We therefore define
@@
which we use to normalize the energy of our system. We caution that $E_{\mathrm{tot}}$ is not precisely invariant due to boundary effects and truncation error from evolving $\tau$ instead of $\alpha(T^{0}_{0}+\rho u^{0})$, leading to a maximum deviation of ${\sim}0.5\%$ in the total energy over the course of the evolution. However, we have confirmed that our results are the same if we normalize by $E(t=0)$ or $E(t)$, indicating that these conservation violations are insignificant for our analysis.
@@
Fig. 2 corroborates this picture; for H1e17_5, $E_{K}$ peaks at a lower energy and begins to decay sooner, indicating that this case is also stable. The magnetic energy decays as well, albeit more slowly than the kinetic energy. If we consider a volume $V=A\ell$, where $A$ is some changing cross-sectional area and $\ell$ is an approximately fixed length, flux conservation implies that $2E_{B}\approx B^{2}V\propto A^{-1}$, so we attribute this slow decay in magnetic energy to the slow diffusion or expansion of the flux tube as demonstrated in Fig. 3.
@@
For H7e17_5, which has the strongest magnetic field, the flux tube successfully emerges. This emergence process is shown in Figs. 4 and 5. The flux tube stretches out into a characteristic umbrella shape due to the decreasing scale height as it rises, similar to what occurs in solar studies (e.g., Fig. 4 of Ref. Toriumi and Yokoyama (2011)). As it reaches the surface, the flux tube rapidly expands and pushes out a wave of material in front of it (see Fig. 6).
@@
Figure 6: A comparison of $|B|$ (left) and $\rho$ (right) in the $r$-$\phi$ plane from H7e17_5 at $t=0.34~\textrm{ms}$. Much of the flux tube remains near the surface, but the wave of material pushed out by it is also visible. The dashed black line marks the location of the stellar surface from the TOV solution, and the dot-dashed green lines mark the boundaries of the flux tube at $t=0$.
To measure the emergence time and properties of the entrained material, we integrate the mass flux $\dot{M}$ across surfaces at $r=8.35~G\textrm{M}_{\odot}c^{-2}\approx 12.33~\mathrm{km}$, $r=8.4~G\textrm{M}_{\odot}c^{-2}\approx 12.40~\mathrm{km}$, $r=8.5~G\textrm{M}_{\odot}c^{-2}\approx 12.55~\mathrm{km}$, $r=8.6~G\textrm{M}_{\odot}c^{-2}\approx 12.70~\mathrm{km}$, and $r=8.7~G\textrm{M}_{\odot}c^{-2}\approx 12.85~\mathrm{km}$. We define the emergence time as the peak mass flux $\dot{M}$ across the surface at $r=8.35~G\textrm{M}_{\odot}c^{-2}$, which is the approximate location of the remnant’s surface. By looking for where $\dot{M}$ peaks in the other integration surfaces, we can measure the approximate speed of the flux tube at emergence, which we do by fitting a quadratic polynomial for $r(t)$ to these peak times and taking its time derivative.
Fig. 7 shows the time-integrated mass across each integration surface along with the estimated emergence time. The flux tube emerges at $t_{\mathrm{em}}\approx 0.216~\textrm{ms}$ for H7e17_5 with a speed of $v_{\mathrm{em}}\approx 0.038c$. The total entrained mass peaks around $M\approx 1.5\times 10^{-3}~\textrm{M}_{\odot}$ shortly after emergence, but it rapidly declines as material falls back onto the remnant. The mass across each integration surface also rapidly decreases with distance, suggesting that little to no mass is unbound [^1] A better way to estimate unbound ejecta is using the geodesic and/or Bernoulli criteria, but this must be done farther away from the remnant to be accurate. Because our simulations only span $\pi/9,\pi/9]$ in the $\theta$ direction with periodic boundary conditions, the emerging fluid begins to interact with itself not long after passing the surface at $r=8.7~G\textrm{M}_{\odot}c^{-2}$. Therefore, the dynamics of the fluid will not be reliable at a radius where a proper ejecta calculation would be meaningful..
@@
Unsurprisingly, Fig. 2 demonstrates that H7e17_5 has larger $E_{B}$ and $E_{K}$ than in the weak-field runs. $E_{K}$ follows the same qualitative trend as the other runs, with a rapid rise followed by a peak and slow decay in time (possibly with small oscillations). Unlike the other runs, however, $E_{B}$ immediately decays, while it remains nearly constant in H1e17_5 and grows for some time in both H1e16_5 and H6e15_5. There is also a crossover between $E_{B}$ and $E_{K}$ at the same approximate time emergence occurs, though, as will be discussed later, this may simply be coincidence.
@@ III.2 Effects of density perturbation
We now consider the effects of the perturbation by analyzing the non-perturbed cases and the $k=9$ cases, beginning with H1e17_0. Since H1e17_5, which has a density perturbation, does not emerge, it is not surprising that H1e17_0 does not, either, and instead rises only for a short time before stalling. As shown in Fig. 8, $E_{K}$ peaks at ${\sim}1/3$ its value in H1e17_5, and it rapidly decays by nearly an order of magnitude in ${\sim}0.3~\textrm{ms}$. Additionally, though the evolution of $E_{B}$ in H1e17_0 is initially similar to H1e17_5, it does not experience the same slow rise, but instead slowly decays over time.
Figure 8: Like Fig. 2, but comparing the cases with different perturbations ($A=0$, $k=9$, and $k=18$). The dotted vertical lines mark the emergence times for H7e17_0, H7e17_5, and H7e17_5_k9.
Comparing the H7e17_0 case to H7e17_5, we find that the perturbation has very little effect on the dynamics of the system. This is consistent with our initial conditions, as the change in density from our equilibrium conditions on the flux tube ($\Delta\rho/\rho_{0}\approx 0.6$) is much larger than the change in density from the perturbation ($A=0.05$). According to Fig. 8, the presence of a perturbation does not affect the energetics of the system in any significant way. However, there is a small effect on both the emergence time and the mass pushed out by the emerging flux tube. Fig. 7 shows that emergence in H7e17_0 is slightly delayed to $t_{\mathrm{em}}\approx 0.224$ and slightly slower at $v_{\mathrm{em}}\approx 0.034c$ compared to H7e17_5, though the entrained mass is slightly larger. Nevertheless, the mass again rapidly decays with distance, suggesting that the amount of unbound ejecta is small.
H1e17_5_k9 demonstrates sustained growth in the kinetic energy compared to H1e17_5 due and therefore rises further, which at first glance may suggest that $k=9$ may be closer to the critical wavenumber than $k=18$. However, the flux tube still fails to emerge, and the kinetic energy never exceeds what is observed in Fig. 2 for H0_0_5. At early times, the growth rate is nearly identical for both perturbation modes, so it seems that the differences in evolution are not necessarily because $k=9$ initially excites the Parker instability but cannot overcome the stratification of the system near the surface; rather, it seems that the dynamics of both $k=9$ and $k=18$ are dominated by the hydrodynamic perturbation, and $k=9$ suppresses it less due to reduced magnetic tension.
On the other hand, the H7e17_5_k9 case emerges at $t_{\mathrm{em}}\approx 0.194~\textrm{ms}$ with a velocity of $v_{\mathrm{em}}\approx 0.048c$, somewhat earlier than H7e17_5, and there is evidence that the kinetic energy is marginally larger. When comparing the kinetic energy evolution at early times for perturbed cases to H7e17_0, we find a maximum relative difference of ${\sim}9\%$ for H7e17_5_k9 compared to ${\sim}2\%$ for H7e17_5. This might suggest the $k=9$ perturbation grows faster, but we note that this difference does not demonstrate clear monotonic growth, so we claim with any confidence that this is more than a numerical effect. Additionally, assuming $E_{K}\propto t^{p}$ for some growth scale $p$, we do not see any meaningful difference in $p$ between the different perturbations. Fig. 9 indicates that $d\log E_{K}/d\log t$ is effectively identical for all three cases, even at early times. Therefore, the growth timescale for the Parker instability must be much longer than the emergence timescale. This is not truly surprising, as the Parker instability does not usually dominate over other buoyant instabilities until it becomes nonlinear (see, e.g., Refs. Matsumoto et al. (1993); Nozawa (2005)).
Figure 9: The growth rate $d\log E_{K}/d\log t$ as a function of time for H7e17_0, H7e17_5, and H7e17_5_k9.
@@ III.4 Effects of resolution
There is little consistency in the emergence time, and the effects of surface oscillations make it difficult to determine exactly where emergence occurs, particularly for the LR case. Though the flux tube is well resolved for the LR, MR, and HR cases (and arguably adequately resolved for the VLR case), it is likely that we are not in a completely convergent regime, even if we ignore the measured emergence times.
@@ IV Discussion
For our magnetic field configuration, we observe that only extremely large fields are capable of emerging. Because our field is a single twisted loop surrounding the entirety of the star, however, there is a strong magnetic tension force which, using Eq. 24, initially exceeded $50\%$ of the gravitational force on the flux tube in some of our tests. Alternate field configurations may be less susceptible to this tension force, but numerical simulations suggest that post-merger fields, independent of their initial (realistic) conditions, are dominated by a large-scale toroidal component Aguilera-Miret et al. (2022); Gutiérrez et al. (2025), which limits the possibilities for alternate field structures.
@@
Even if this were not the case, this likely is only important for fields of ${\sim}10^{17}~\textrm{G}$; we observe that the evolution of the flux tube at weaker field strengths around $10^{16}~\textrm{G}$ are dominated by ordinary hydrodynamic buoyancy. Therefore, it seems unlikely that altering the field configuration is enough to help these features to emerge. As stated in Sec. II.1, a rotating model is also unlikely to improve emergence, as rotation tends to inhibit convective processes, not enhance them Acheson and Gibbons (1978); Spruit (1999); Gao et al. (2026).
@@
This result may initially seem at odds with works suggesting that strong magnetically driven outflows may develop from a BNS remnant, such as Most and Quataert (2023) and Combi and Siegel (2023). This is not necessarily the case; the demonstrated flare-like features and outflows in Most and Quataert (2023) originate from strong magnetic fields exceeding $10^{17}~\textrm{G}$. Similarly, the post-merger jet in Combi and Siegel (2023) develops in a remnant with an average toroidal field strength nearing and isolated pockets exceeding $10^{17}~\textrm{G}$. However, we again stress that such field strengths are atypically large compared to observation of magnetars and what is expected in realistic merger remnants. Even under these circumstances, such fields are not a guarantee that magnetically-driven outflows develop from the remnant. Musolino et al. (2025), for example, suggest that, even with an extremely large magnetic field, these outflows may instead originate from the disk.
@@
Similar to alternate configurations, buoyant instabilities like the Parker instability are likely only important for a small range of field strengths in excess of $10^{17}~\textrm{G}$. Despite a strong longitudinal perturbation, the Parker instability does not appear to develop in H1e17_5; rather, the flux tube oscillates. This agrees with results from Refs. Musolino et al. (2025); Jiang et al. (2025), for instance, which suggest that the post-merger disk can be unstable to the Parker instability, but the remnant itself should be stable, even for a very large magnetic field. Larger fields like H7e17_5 may be Parker-unstable, but the emergence time scale is so short that the instability cannot develop quickly enough to be dynamically important (as seen when compared with H7e17_0). There likely exists some middle ground between these two states where the Parker instability develops quickly enough to accelerate the emergence time scale, but the required field strengths may not be achievable in a realistic merger setting.
@@
As we consider cases which do emerge, it is interesting to note that whether or not emergence occurs seems to depend only mildly on resolution so long as the buoyant feature is adequately resolved. Though this is a necessary condition for emergence, one would not necessarily expect it to be sufficient; under-resolving $H_{\rho}$ could lead to an incorrect density difference between the ambient medium and the interior of the buoyant feature, therefore leading to an incorrect emergence time scale. Similarly, under-resolving $H_{P}$ could cause the buoyant feature to expand too much or too little as it rises, thus indirectly affecting the density difference and limiting buoyancy. In our runs, $H_{\rho}\approx 2.43~\mathrm{km}$ and $H_{P}\approx 827~\mathrm{km}$ at the shear layer, but both these quantities will rapidly decrease as the flux tube approaches the surface. Therefore, $H_{\rho}$ will only be moderately resolved at the beginning for the VLR run, and $H_{P}$ will only be resolved by the HR and possibly MR runs to start. Though we see some differences in the emergence time (which may not be entirely accurate for H7e17_5_B), they are small enough that they would be insignificant in the context of a full merger simulation. This suggests that the timescales shown in global BNS runs with $\Delta x\gtrsim 180~\mathrm{m}$ are likely reasonable despite the limited resolution.